### v1

# Thinking Before Running! Efficient Code Generation with Thorough Exploration and Optimal Refinement

Abstract
Code generation is crucial in software engineering for automating the coding process efficiently. While test-time computation methods show promise, they suffer from high latency due to multiple computation rounds. To overcome this, we introduce ThinkCoder, a framework that combines thorough exploration with optimal refinement. The exploration phase diversifies the solution space by searching for potential solutions, followed by a refinement phase that enhances precision. This approach allows us to select the best solution through careful consideration before taking action, avoiding excessive trial and error. To further minimize test-time computation overhead, we introduce preference-driven optimization with Reinforced Self-Training (ReST), which uses exploration trajectories from ThinkCoder to guide LLM’s evolution. By learning preferences, this approach improves LLM’s exploration efficiency, reducing computational costs while maintaining accuracy. ThinkCoder boosts the performance of multiple base LLMs, excelling on benchmarks like HumanEval and MBPP. Compared to SOTA models, it improves Pass@1 by 1.5% over MapCoder with just 21.7% of the computation cost. Against AgentCoder, ThinkCoder achieves a 0.6% higher Pass@1 after 2 rounds, outperforming AgentCoder’s 5 rounds. Additionally, ReST with success trajectories enhances efficiency, allowing models like LLaMA2-7B to achieve competitive results using only 20% of the computational resources. These results highlight the framework’s effectiveness and scalability. The code will be released upon acceptance.
Thinking Before Running! Efficient Code Generation with
Thorough Exploration and Optimal Refinement
Xiaoqing Zhang1   Yuhan  Liu1   Flood Sung2  Xiuying Chen311footnotemark: 1     Rui  Yan1thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China  2MoonshotAI 3Mohamed bin Zayed University of Artificial Intelligence {xiaoqingz, yuhan.liu, ruiyan}@ruc.edu.cn {floodsung}@moonshot.cn {xy-chen}@pku.edu.cn
1 Introduction
Recent research advances indicate that large language models (LLMs) have demonstrated remarkable capabilities in various programming-related domains, such as code generation Zheng et al. (2023); Chaudhary (2023); Dong et al. (2024); Li et al. (2024a), code refinement Chen et al. (2023); Guo et al. (2024); Zheng et al. (2024); Ridnik et al. (2024); Liu et al. (2024); Zhang et al. (2024), and software testing Li et al. (2023); Jalil et al. (2023); Wang et al. (2024); Jones et al. (2024). By utilizing extensive training data, LLMs can understand complex programming tasks, generate syntactically accurate code, and even enhance the quality of their solutions through iterative improvement.
In the context of code generation, researchers have made significant progress in improving prompts to encourage LLMs to generate higher-quality answers. This includes two main directions: optimizing the path for answer exploration and incorporating rich reflection information. In the first direction of path optimization, for example, Self-Planning Jiang et al. (2024) reduces problem complexity in code generation by introducing a planning phase, which refines the process of solution generation. Similarly, CodeCoT Huang et al. (2023a) improves the exploration by employing a Chain-of-Thought prompt, enabling models to reason through each step systematically. In the second direction, the integration of reflection, test cases, and feedback is progressively being used to steer and enhance the generation process. For instance, RethinkMCTS Li et al. (2024b) integrates verbal feedback into its Monte Carlo Tree Search framework to correct errors and improve code generation. Similarly, AgentCoder Huang et al. (2023b) boosts efficiency and accuracy by using interactions between multiple agents to provide valuable input.
Despite their impressive capabilities, a critical challenge remains in fully exploiting the potential of LLMs for high-quality code generation, as their performance is still constrained by issues such as high computation costs and limited effectiveness in integrating iterative feedback. To enhance efficiency and minimize trial-and-error costs, it’s essential to define the improvement direction clearly before implementing refinement. While existing methods like MapCoder, which leverages four LLM-based agents, consumes a considerable amount of computational resources.
To address these challenges, we propose ThinkCoder, a novel framework for efficient code generation that combines thorough exploration and optimal refinement. First, we simplified our multi-agent framework, which now consists only of the Exploration Agent and the Execution Agent. The Exploration Agent is responsible for deep thinking and expanding exploration with various codes and tests. The Execution Agent independently verifies code during the optimal refinement process, without relying on LLMs. By utilizing a dynamically created testing pool, ThinkCoder ensures accurate verification of the generated code and offers reliable guidance for further optimization. Second, to tackle the additional computational load resulting from extensive exploration, we incorporate a preference-driven optimization phase using Reinforced Self-Training (ReST) in ThinkCoder. ReST leverages successful exploration trajectories to train LLM, enabling it to generate optimal solutions more efficiently during exploration. This approach significantly reduces the computational cost of test-time computations, enabling models like LLaMA2-7B to achieve notable improvements in both performance and efficiency.
Our contributions are as follows:
$\bullet$ We propose ThinkCoder, an efficiency framework for code generation that incorporates only two agents: an exploration agent and an execution agent that enables thorough exploration and optimal refinement, ensuring high-quality code generation.
$\bullet$ We introduce preference-driven optimization with ReST, using exploration trajectories from ThinkCoder to enhance the generation and verification capabilities of LLMs, further reducing computational overhead during testing.
$\bullet$ We have demonstrated the effectiveness of ThinkCoder on multiple benchmark datasets and across different scales of LLMs, achieving state-of-the-art performance in code generation while optimizing computational efficiency.
2 Related Work
Test-time Compute. Test-time computing methods use self-play and additional verifiers to enhance performance, despite the computational overhead. For example, training a Verifier with Preference Reward Model data and combining strategies like Best-of-N Weighted, Beam Search, and Lookahead Search can help Snell et al. (2024). ToolLLM Qin et al. (2023) evaluates tool usage by comparing success rates and solution quality to ChatGPT-ReACT, employing ToolEval as a Verifier for optimal API calls using depth-first search. However, API-based Verifiers are limited by training data and struggle with generalization. In contrast, our approach employs a task-based testing tool as a Verifier, offering a simpler and more effective solution that is closely aligned with the specific task requirements.
Multi-Agent Code Generation. Deriving multiple agents by mimicking the different stages of human programming can effectively improve both the quality and efficiency of code generation. For instance, AgentCoder Huang et al. (2023b) introduces programmer agents, test design agents, and test execution agents, integrating code generation, test case generation, and test feedback into the code generation process. Similarly, MapCoder Islam et al. (2024) introduces retrieval agents, planning agents, coding agents, and debugging agents, iterating on code generation quality through more frequent agent communication and reflection. However, MapCoder incurs higher computational overhead and relies solely on sample I/O for code execution, limiting test case coverage and robustness. In contrast, our framework, ThinkCoder, addresses these limitations by self-verification efficiently without excessive computational costs.
Figure 1: The end-to-end process of ThinkCoder involves $k$ thorough exploration steps followed by $n$ optimal refinement cycles. The Exploration Agent generates $k$ codes and tests simultaneously, storing results in a Testing Pool. The optimal refinement phase includes self-verification that selects the optimal solution with the Execution Agent and aggregate reflection for the instruction update. The optimal refinement will be repeated recursively for $n$ cycles, ultimately leading to the final solution.
Instruction Tuning for Code. High-quality code generation models often require instruction fine-tuning based on large-scale pre-trained language models. The fine-tuning data generally comes from real-world collections or is artificially synthesized. For example, Code Alpaca Chaudhary (2023) combines the SELF-INSTRUCT strategy with code fine-tuning, CODELLAMA Roziere et al. (2023) generates questions, solutions, and test cases through prompts, and OctoPack Muennighoff et al. (2023) directly collects data from Git. LLMs fine-tuned with instructions possess stronger code generation capabilities. However, existing methods often lack comprehensive guidelines for data collection, volume management, and fine-tuning strategies. To address this gap, we present a detailed fine-tuning strategy for Llama2-7B using instruction-tuning data collected via CodeQwen Bai et al. (2023), providing a transparent and reproducible methodology that supports the advancement of instruction-tuned code generation models.
3 ThinkCoder
Figure 2: Trajectory collection with ThinkCoder and its application in ReST training for LLMs. We collect success trajectories offline based on the verification with ground truth tests, ensuring the solution aligns with human preferences. For reflection, we use LLM-generated tests, ensuring LLMs specifically address their own mistakes.
3.1 Overall Architecture
Figure 1 illustrates the structure of ThinkCoder. Given an instruction $I=[p,\text{prompt}]$ for problem $p$, where prompt is the task description for code generation, the Exploration Agent produces $k$ candidate solutions. Each solution is paired with $m$ test cases, and the temperature $t$ of the LLM’s generation is adjusted to promote diversity. These test cases are then aggregated into a Testing Pool $TP$. We define the exploration results as $E=[G,TP]$, where $G=\{g_{i}\mid g_{i}=\text{LLM}(p,t)\text{ for }i\in\{1,2,\dots,k\}\}$, and $TP\in\mathbb{T}^{k\times m}$, where $\mathbb{T}$ represents each test. The Execution Agent sequentially executes each candidate solution in $G$ using the test cases from $TP$ to compute pass rates. The code $g_{s}$ with the highest pass rate is selected for further refinement. Here we define the pass rate of $g_{i}$ as $r_{g_{i}}=\frac{\text{number of tests passed by }g_{i}}{|TP|}$ and $g_{s}=\arg\max_{i}r_{g_{i}}$. Feedback $f$ from failed test case $f_{t}$, along with the current code $g_{s}$, generated test cases $tests$, and problem $p$, is used to update the instructions $I=[p,g_{s},f_{t},f,tests,prompt]$ for the next exploration. This iterative process allows the Exploration Agent to generate improved solutions and test cases over $n$ iterations. To avoid unnecessary computation on well-solved problems, if $r_{g_{i}}$ is greater than $\theta$, the refinement process is terminated, and the code $g_{s}$ is returned as the final solution. Please refer to the ThinkCoder algorithm 1 for more detailed information on these operations. Below, we detail the implementation of Thorough Exploration, Optimal Refinement, Exploration&Execution Agent, and Testing Pool.
Algorithm 1 ThinkCoder
1:Initialize the testing pool TP=[], set the globally max pass rate $g_{p}=0$ and best solution $g_{s}=None$
2:Set the hyperparameters $t=0.5,k=5,n=5,m=3,\theta=0.8$
3:Set the instruction $prompt$ for code generation task and instruction for task $p$ as $I=[p,prompt]$
4:for each $iter\in n$ do
5:     Initialize the locally max pass rate $l_{p}=0$ and best solution ${l_{s}}={None}$ for iteration $iter$
6:     Randomly select $m$ successful test cases from TP as $tests$
7:     Select a new type of error feedback for test $f_{t}$ from TP as $f$
8:     if $f_{t}$ and $f$ is not None then
9:         Update the instruction $I=[p,g_{s},f_{t},f,tests,prompt]$
10:     end if
11:     Get code solutions $G^{\prime}$ and test cases $TP^{\prime}$ from exploration agent for the instruction $I$ within $k$ iterations
12:     Deduplicate testing pool TP’ by AST
13:     for each $i\in|G^{\prime}|$ do
14:         Get the pass rate $r_{g_{i}}$ for solution $g_{i}$ from execution agent with the input $TP^{\prime}$
15:         if $r_{g_{i}}>l_{p}$ then
16:              Set $l_{p}=r_{g_{i}},l_{s}={g_{i}}$
17:         end if
18:     end for
19:     if $l_{p}>g_{p}$ then
20:         Set $g_{p}=l_{p},g_{s}=l_{s}$
21:         Add TP’ to TP
22:     end if
23:     if $g_{p}>\theta$ then
24:         return $g_{s}$ as the best solution
25:         Terminate optimal refinement iteration
26:     end if
27:end for
3.2 Thorough Exploration
Before running the generated code and performing corrections, ThinkCoder conducted extensive thinking during the Thorough Exploration phase. By generating $k$ answers at once, it expanded the search area. The search for multiple answers was achieved by increasing the LLM’s temperature. With the increasing temperature, the probability distribution of the LLM becomes flatter, reducing the differences between word probabilities and increasing the entropy of the distribution, which enhances the diversity of the generated text. The first iteration of exploration aims to generate diverse code and comprehensive test cases, including regular tests, boundary tests, and performance tests. While the subsequent iterations, focus on the exploration of code correction for specific errors. The exploration is carried out by the Exploration Agent, which is the only LLM used in ThinkCoder. The prompt design refers to Appendix A.3.
3.3 Optimal Refinement
To solve the problem efficiently, it is essential to first identify the optimal solution through careful consideration before making any code corrections. Therefore, Optimal Refinement involves selecting the best solution via self-verification and gathering the necessary information for deeper exploration and improvement in the next exploration. During self-verification, code quality is evaluated by the pass rate, determined by the Execution Agent running candidate code against all test cases. The code with the highest pass rate is selected for further refinement. From this, we randomly select an unresolved error to target for the current round of improvements. This ensures all errors are eventually addressed. The refinement instruction is created using the following information: the problem, the best solution identified so far, feedback, and test cases. After each refinement, the solution’s pass rate is evaluated. If it surpasses any previous solution, it is marked as the best. Through multiple iterations, the best solution is continuously corrected and optimized, leading to a gradual increase in the pass rate within the testing pool. For more details, please refer to the Appendix A.3.
3.4 Exploration & Execution Agent
ThinkAgent consists of two agents: the exploration agent and the execution agent. The exploration agent uses LLMs to search for diversity code and test cases, while the execution agent runs the provided code snippets and test cases in a local environment to generate feedback. To improve the efficiency of the execution agent, we leverage multiprocessing to validate the code in parallel.
3.5 Testing Pool
The Testing Pool contains the test cases generated by the exploration agent which contains significantly more test cases than the ground truth tests. These test cases are categorized into three types: regular testing, boundary testing, and performance testing. A higher pass rate in the testing pool reflects stronger code robustness, making it essential to ensure the diversity and accuracy of the test cases. Several key operations contribute to achieving this. Abstract Syntax Tree (AST)  Backus et al. (1960) is applied to identify and remove duplicate test cases in the testing pool, eliminating unnecessary time costs associated with redundant tests. Additionally, after each iteration, the testing pool only selects test cases that can improve the code’s pass rate, thus enhancing the diversity and accuracy of the pool. Finally, the testing pool will select test cases with distinct errors and use feedback to update the code’s improvement goals, driving continuous optimization.
[…]
LLama2-7B-Chat 0 27.8 22.9 12.9 11.2 18.7
1 35.3(27.0%) 29.0(26.6%) 15.9(23.3%) 11.6(3.6%) 23.0(23.0%)
2 37.5(34.9%) 29.5(28.8%) 17.1(32.6%) 12.2(8.9%) 24.1(28.9%)
5 38.1(37.1%) 29.8(30.1%) 21.3(65.1%) 12.9(15.2%) 25.5(35.4%)
CodeQwen1.5-7B-Chat 0 73.1 67.8 79.5 71.0 72.9
1 80.2(9.7%) 74.7(10.2%) 90.9(14.3%) 80.5(13.4%) 81.6(11.9%)
2 84.4(15.5%) 75.0(10.6%) 92.1(15.8%) 80.5(13.4%) 83.0(13.9%)
5 86.0(17.6%) 75.0(10.6%) 93.3(17.4%) 80.5(13.4%) 83.7(14.8%)
Kimi 0 70.0 67.4 81.6 73.4 73.1
1 77.2(10.3%) 73.5(9.1%) 87.2(6.9%) 79.9(8.9%) 79.5(8.8%)
2 78.7(12.4%) 74.5(10.5%) 87.2(6.9%) 79.3(8.4%) 79.9(9.3%)
5 79.5(13.6%) 74.8(11.0%) 86.6(6.1%) 79.3(8.4%) 80.1(9.6%)
GPT-4-Turbo 0 81.2 74.4 85.7 81.0 80.6
1 89.5(5.1%) 78.2(5.1%) 90.2(5.3%) 84.8(4.7%) 85.7(6.3%)
2 89.5(5.1%) 79.9(7.4%) 90.8(6.0%) 84.8(4.7%) 86.3(7.1%)
5 89.5(5.1%) 80.0(7.5%) 91.5(6.8%) 84.8(4.7%) 86.5(7.3%)
GPT-4o 0 84.9 78.2 92.5 85.8 85.4
1 88.5(4.2%) 79.9(2.2%) 96.3(4.1%) 87.8(2.3%) 88.1(3.2%)
2 89.6(5.5%) 79.2(1.3%) 97.0(4.9%) 89.0(3.7%) 88.7(3.9%)
5 90.0(6.0%) 79.2(1.3%) 97.0(4.9%) 89.6(4.4%) 89.0(4.2%)
Table 1: ThinkCoder performance of various LLMs. $n=0$ indicates results without the framework. $n>0$ indicates results after $n$ iterations of ThinkCoder, with each exploration generating $k=5$ solutions for optimal refinement at temperature $t=0.5$. Percentages in parentheses show improvement over the base LLMs.
4 Preference-driven Optimization
In ThinkCoder, the generated code evolves as the pass rate on the testing pool improves. Each increase in the pass rate signifies the development of better codes. This progression demonstrates that the exploration agent is gradually producing higher-quality outputs. By collecting the trajectory of these improvements, we aim to fine-tune the model, enabling the LLM to generate superior code and test cases more effectively. This approach helps to refine the exploration process, clarify its direction, and reduce the time required for exploration. Figure 2 describes the process of trajectory collection and fine-tuning of the LLM. We will explain this process from the perspectives of Data Collection, the Model Training and Inference.
4.1 Data Collection
The process of data collection for LLM fine-tuning involves two key steps. First, an exploration agent, denoted as $\mathcal{M}_{0}$, generates diverse low-probability data $D_{M_{0}}$ by sampling at various temperatures. Second, this data is further refined using reflection to produce an enhanced dataset $R_{M_{0}}$. In our data generation task, offline data collection starts with a training set $D=\{x_{i}\}_{i=1}^{N}$, where $N$ is the number of instances in the dataset. We generate high-quality preference data $D_{M_{0}}\cup R_{M_{0}}$ for LLM fine-tuning through the temperature-based and reflection-based generation process.
Temperature-based Generation. During temperature-based generation, we obtain $D_{M_{0}}$ to enhance the diversity of responses. For a given problem $x$, the language model $\mathcal{M}_{0}$ is used to explore $k$ times, generating $\{y^{j}\}_{j=1}^{k}$, where $y^{j}\sim\mathcal{M}_{0}(y|x)$. During exploration, a relatively large temperature value $t$ is set for $\mathcal{M}_{0}$. The generated outputs are then evaluated using the ExecuteAgent, which computes $E(x,y^{j})=(pass\_rate,feedback)$. For any sample where $pass\_rate==1$, the corresponding instance $(x,y^{j})$ is added to the training dataset $D_{M_{0}}$.
Reflection-based Generation. The primary purpose of generating $R_{M_{0}}$ is to enhance the ability of error correction. For a given sample $x$ and the feedback $E(x,y^{j})=(pass\_rate,feedback)$ from the ExecuteAgent, for samples with $pass\_rate=0$, we add the feedback to $x$ and use the language model $\mathcal{M}_{0}$ to perform temperature-based exploration on the input $x^{\prime}=(x,feedback)$ and get the output $y^{\prime j}$. Then we also use the execution agent to obtain the right correction answer with $E(x^{\prime},y^{\prime j})=(pass\_rate^{\prime},feedback^{\prime})$. Finally, we add the instance $(x^{\prime},y^{\prime j})$ with the $pass\_rate^{\prime}==1$ to the reflection-based trajectories $R_{M_{0}}$.
4.2 Model Training and Inference
Generating high-probability text does not necessarily align well with human preferences, which may lead to biases in various tasks Perez et al. (2022). Reinforced Self-Training(ReST) mitigates this issue effectively by collecting offline data aligned with human preferences for online training Gulcehre et al. (2023). We train the base model $\mathcal{M}$ with the success trajectory dataset $D_{M_{0}}\cup R_{M_{0}}$:
[…]
In the inference process, two distinct configurations are used. The first configuration utilizes the fine-tuned base model, where the input $x$ is fed directly into the model to generate text with a temperature setting of $t=0$. This setup is designed to assess whether the data collected through ThinkCoder enhances the quality of the generated output. The second configuration regards the fine-tuned base model as the exploration agent within the ThinkCoder framework. This approach generates and refines solutions using $k$ explorations and $n$ refinements, with various temperature settings $t$. It aims to investigate the framework’s ability to improve the generated solutions through the success trajectories and the ThinkCoder framework. Comprehensive experimental results are presented in the evaluation section.
5 Experiments
5.1 Dataset
We utilized the MBPP Austin et al. (2021) and HumanEval Chen et al. (2021) datasets to evaluate the effectiveness of our approach. To ensure a comprehensive assessment, we tested their enhanced versions—MBPP-ET and HumanEval-ET Dong et al. (2023)—which include 80 and 35 times more test cases than the original datasets, respectively. Regarding the test data, the original problem set sizes for MBPP and HumanEval are 257 and 164, while their extended versions contain 378 and 164 problems, respectively. For training, we started with 374 MBPP training examples and 90 validation examples, combining them as the initial dataset. Using temperature-based and reflection-based generation methods, we expanded this set to over 2,000 examples for ReST training.
5.2 Baselines
In terms of LLM base model selection for the application of ThinkCoder, we considered a range of advanced instruction-tuned models, including the general-purpose models LLama2-7B-Chat, Kimi [^1] https://platform.moonshot.cn, GPT-4-Turbo(gpt-4-1106-preview) and GPT-4o, as well as the code-specific model CodeQwen1.5-7B-Chat. We conducted a comprehensive evaluation of ThinkCoder’s Pass@1 performance on MBPP and HumanEval across LLMs of different parameter sizes. For the test-time optimization process, we used the base model LLama2-7B to evaluate performance changes on the MBPP dataset. Comparisons were made between results obtained before and after fine-tuning, using either the original training data or the success trajectories, and incorporating the ThinkCoder framework or excluding it. For the composite model, we compared it with the current State-of-the-Art (SOTA) models, Self-Collaboration Dong et al. (2024), MapCoder Islam et al. (2024), and AgentCoder Huang et al. (2023b).
Models Optimal Refinement Iterations
0 1 2 3 4 5
LLama2-7B 14.7 22.7 24.9 27.7 32.7 36.9
LLama2-7B(SFT) 27.8 35.3 37.5 37.6 38.0 38.1
LLama2-7B(ReST) 32.3 46.2 44.8 46.1 45.7 46.1
Table 2: Performance comparison of the LLama2-7B model under different settings: baseline without fine-tuning, supervised fine-tuning (SFT) with original train dataset, and Reinforced Self-Training (ReST) with success trajectories from ThinkCoder on the MBPP dataset.
Figure 3: (a) The Pass@1 metric of the baseline models under different temperature $t$. (b) The Pass@k metric of the baseline models, where $k$ represents the exploration budget. (c) The variation of the Pass@1 metric for each baseline model under the ThinkCoder framework as the optimal refinement budget $n$ increases.
5.3 Implementation details
We set different parameter combinations for different experimental settings. To validate the improvement of the base LLMs, we set the exploration budget $k=5$, the temperature received by the exploration agent $t=0.5$, the number of optimal refinements $n=5$, and the budget control threshold $\theta=0.8$. The exploration agent generates $m=3$ test cases per code. To compare ThinkCoder with state-of-the-art (SOTA) results, we set the hyperparameters $k$, $t$, $n$, $\theta$, and $m$ to 20, 1.0, 2, 1.0, and 3, respectively. In the data collection phase, the exploration agent $\mathcal{M}_{0}$ is specified as CodeQwen1.5-7B-Chat because it is free and performs well. During ReST training, we set the learning rate to $2\times 10^{-5}$ and train for 2 epochs. We employ Low-Rank Adaptation (LoRA) Hu et al. (2021) to train LLMs.
Figure 4: The variation in Pass@1 performance of the exploration code and test cases on the MBPP dataset under ThinkCoder, as evaluated by CodeQwen1.5-7B-Chat as the base LLM.
6 Evaluation
Table 1 illustrates the performance changes of various base LLMs on the MBPP and HumanEval datasets as the Optimal Refinement iteration $n$ increases. Table 3 compares the performance of ThinkCoder and SOTA models, using GPT-4-Turbo as the base LLM. Table 2 presents the performance changes of LLama2-7B before and after fine-tuning with trajectory data collected by ThinkCoder. Figure 3 shows the impact of changes in the hyperparameter temperature $t$, exploration budget $k$, and refinement iterations $n$ on the Pass@1 and Pass@k metric. Figure 4 depicts the trends in the accuracy of generated test cases and generated codes as the optimal refinement iterations in the ThinkCoder increase. Table 4 compares ThinkCoder’s computation load with MapCoder at similar settings.
[…]
Self-Collaboration 90.2 70.7 78.9 62.1 75.5
MapCoder 83.1 57.7 93.9 82.9 79.4
AgentCoder 91.4 91.4 89.6 76.2 87.2
ThinkCoder 90.5 80.5 93.9 86.0 87.7
Table 3: End-to-end results of ThinkCoder and state-of-the-art (SOTA) baselines.
6.1 Performance on Code Generation
Table 1 shows that ThinkCoder achieves significant improvements in scenarios with fewer standard test cases, with performance on MBPP and HumanEval improving by 4%-70% after 5 iterations of refinement. In contrast, gains in more extensive scenarios, like MBPP-ET and HumanEval-ET, are more modest, ranging from 1%-35%. ThinkCoder’s generated test cases effectively verify code, reducing the need for manual tests, but its performance saturates in complex scenarios, making it better suited for simpler tasks. Table 3 further demonstrates the potential of ThinkCoder. When the number of exploration iterations $k$ is extended to 20, ThinkCoder surpasses AgentCoder’s performance, which needs 5 refinement rounds, in just 2 iterations.
Models Exploration k HumanEval MBPP
Pass@1 Refinement n Tokens (k) Pass@1 Refinement n Tokens (k)
MapCoder 5 93.9 5 10.41 83.1 3 4.96
ThinkCoder 5 90.2 1 1.92 89.5 1 1.44
Table 4: The token usage for agent responses and the required refinement iterations in ThinkCoder for HumanEval and MBPP, compared to MapCoder (as ThinkCoder and MapCoder share similar configurations), with GPT-4-Turbo as the base LLM. The tiktoken package is used to calculate the response token usage.
6.2 Performance with ReST Training
From the results presented in Table 2, it is clear that ReST fine-tuning on LLama2-7B using successful trajectories collected by ThinkCoder significantly enhances the model’s core capabilities. Firstly, LLama2-7B(ReST) demonstrates superior performance compared to both LLama2-7B(SFT) and LLama2-7B at iteration 0, highlighting that ReST training improves the exploration agent’s ability to identify better solutions. Secondly, with a temperature setting of 0.5, the model achieves optimal performance with only a single round of refinement, underscoring the efficiency of ReST fine-tuning in reducing computational overhead.
6.3 Evolution of Generated Code and Tests
After each round of optimal refinement, we evaluated the generated solutions and testing pools. Using the ground truth test set, we analyzed the performance changes of the solutions in each round. Meanwhile, the ground truth code was used to assess the changes in the testing pools. As shown in Figure 4, the accuracy of both the solutions and testing pools improved simultaneously. This demonstrates that ThinkCoder is capable of generating continuously optimized test cases while also refining the quality of the solutions.
6.4 Hyperparameters study
We conducted hyperparameter selection experiments on the MBPP dataset. As shown in Figure 3(a), models exhibit optimal performance at different temperature settings. LLama2-7B-Chat, CodeQwen1.5-7B-Chat and Kimi improve in Pass@1 as temperature increases, while GPT-4-Turbo and GPT-4o perform better at lower temperatures. As depicted in Figure 3(b), the Pass@k performance of various base LLMs improves with an increasing exploration budget $k$, highlighting the models’ ability to find optimal solutions. However, the performance eventually levels off, suggesting that the Pass@k performance of these models will eventually stabilize. As shown in Figure 3(c), Pass@1 performance of base LLMs improves with increasing refinement iterations $n$, following a trend similar to the budget curve during exploration. However, exploration consistently achieves greater performance gains, underscoring the importance of optimizing the LLM’s answer selection ability for better code quality. For the hyperparameter study of $\theta$ and $m$, please refer to Appendix A.1.
6.5 Cost Analysis
In Figure 4, we compare the computational overhead between MapCoder and ThinkCoder. On the HumanEval and MBPP datasets, a single refinement with ThinkCoder outperforms the result of five iterations of refinement with MapCoder, while using fewer tokens. This is because MapCoder utilizes four agents that are all powered by LLMs, whereas our framework uses a single LLM to handle both code and test generation, while debugging is directly based on feedback from the Python environment, significantly reducing token usage.
7 Conclusion
In this paper, we introduce ThinkCoder, a framework designed to improve the code generation capabilities of LLMs. The proposed approach incorporates an extensive exploration to generate diverse solutions, followed by an optimal refinement stage that integrates the solutions, tests, and feedback as reflections to enhance performance. To minimize computational costs during exploration, we capture successful trajectories through ThinkCoder, which are then leveraged to fine-tune the LLM. This process not only boosts the model’s performance in code generation but also reduces the number of refinement iterations needed to achieve optimal solutions. We evaluate our framework across LLMs of varying sizes and demonstrate its superior performance over current composite models.
Limitations
In our current approach to managing the testing pool, we primarily focus on selective expansion but lack effective strategies for cleaning noise from the test set. Moreover, the diversity of the test set heavily depends on the generation capabilities of LLMs. Our work highlights the importance of tests and identifies the gap between generated and annotated cases, so we plan to address the enhancement of the testing pool in future work. Furthermore, while ReST training has demonstrated excellent results in enhancing the code generation capabilities of foundational LLMs, validation in ThinkCoder reveals rapid convergence. This indicates that fine-tuning teaches the LLM to find good answers, but the improvement in exploring the overall solution space remains limited. To address this, we look forward to introducing on-policy training methods in future work, combining data collection with fine-tuning to enable the model to evolve continuously during training, generating more diverse answers and further boosting the LLM’s online exploration capabilities.
Appendix A Appendix
A.1 Exploration Budget Control
Figure 5: The relationship between the computational cost and performance of ThinkCoder at different budget control thresholds, where ‘Pass@1’ indicates the performance on the MBPP dataset, and ‘$\theta$’ refers to the budget control threshold that allows the task to execute the next exploration process. ‘Exploration Budget’ represents the ratio of total requests during each iteration.
Figure 5 outlines the number of tasks for the exploration agent per iteration and the overall Pass@1 metric under different budget control thresholds. By increasing the exploration budget control threshold $\theta$, the number of tasks requiring correction in each iteration is significantly reduced, leading to a notable decrease in ThinkCoder’s runtime. However, this approach risks leaving some incorrectly validated tasks uncorrected, which could negatively impact performance. In our experiments, we set $\theta=0.8$, as this value not only progressively improves Pass@1 but also reduces the number of tasks per iteration by nearly half.
A.2 Exploration for Tests
Metrics Tests m/Exploration
1 3 5 10
Pass@1 80.0 81.7 81.7 86.7
Overload(s) 28.7 86.0 112.3 364.5
Table 5: The variation in the Pass@1 metric and the execution agent’s overload is influenced by changes in the number of test cases $m$ during each exploration.
We randomly selected 50 tasks from the MBPP test data to observe the performance improvement and code execution overhead as the number of test cases $m$ increases during each exploration. Table 5 illustrates this trend. The base model we used is Kimi. We observed that increasing $m$ leads to improved Pass@1 performance and results in longer execution times due to the additional effort required for self-verification. To balance time efficiency and performance, we set $m$ for all experiments.
A.3 Prompts for Exploration Agent
Figure 6: ThinkCoder exploration agent prompt example.
Figure 6 illustrates the input received by the exploration agent, including the problem, the best solution, the test cases from the ground truth, the generated test cases from testing pooling, and feedback on test failures.
A.4 Prompts for Execution Agent
Figure 7: ThinkCoder execution agent prompt example.
Figure 7 shows the input of the executable Python environment, including the code and all test cases from the testing pool.
A.5 Example of Thorough Exploration
Figure 8: ThinkCoder exploration examples.
Figure 8 illustrates multiple solutions explored by the exploration agent.
A.6 Example of Optimal Refinement
Figure 9: ThinkCoder refinement examples.
To better understand the process of optimal refinement, Figure 9 illustrates a single refinement iteration. Before refinement, we leveraged the best solution output from the previous iteration, along with failed tests and unmet feedback sampled from the testing pool. Using this reflection information, the solution was updated and revalidated on the test cases in the testing pool. The success rate improved from 0.217 to 0.356, leading us to select the current solution as the best solution.
A.7 Example of Testing Pool
Figure 10: ThinkCoder testing pool examples.
Figure 10 illustrates the test cases in the testing pool generated through multiple explorations by ThinkCoder.

### v2

@@
 To further minimize test-time computation overhead, we introduce preference-driven optimization with Reinforced Self-Training (ReST), which uses exploration trajectories from ThinkCoder to guide LLM’s evolution.
-By learning preferences, this approach improves LLM’s exploration efficiency, reducing computational costs while maintaining accuracy.
-ThinkCoder boosts the performance of multiple base LLMs, excelling on benchmarks like HumanEval and MBPP.
-Compared to SOTA models, it improves Pass@1 by 1.5% over MapCoder with just 21.7% of the computation cost.
-Against AgentCoder, ThinkCoder achieves a 0.6% higher Pass@1 after 2 rounds, outperforming AgentCoder’s 5 rounds.
+This approach enhances LLM’s exploration efficiency via preference learning, cutting costs while maintaining accuracy.
+ThinkCoder boosts the performance with a single LLM, excelling on benchmarks like HumanEval and MBPP.
+Compared to SOTA models, it improves Pass@1 by 3.0% over MapCoder with just 6.4% of the computation cost.
+Against AgentCoder, ThinkCoder achieves a 0.5% higher Pass@1 after 2 rounds, outperforming AgentCoder’s 5 rounds.
 Additionally, ReST with success trajectories enhances efficiency, allowing models like LLaMA2-7B to achieve competitive results using only 20% of the computational resources.
 These results highlight the framework’s effectiveness and scalability.
-The code will be released upon acceptance.
+[^1] The code is available at https://github.com/xiaoqzhwhu/ThinkCoder
 Thinking Before Running! Efficient Code Generation with
 Thorough Exploration and Optimal Refinement
-Xiaoqing Zhang1   Yuhan  Liu1   Flood Sung2  Xiuying Chen311footnotemark: 1     Rui  Yan1thanks:   Corresponding authors.
-1Gaoling School of Artificial Intelligence, Renmin University of China  2MoonshotAI 3Mohamed bin Zayed University of Artificial Intelligence {xiaoqingz, yuhan.liu, ruiyan}@ruc.edu.cn {floodsung}@moonshot.cn {xy-chen}@pku.edu.cn
+Xiaoqing Zhang1,2thanks: This work was done during the internship at Moonshot AI.
+Yuhan  Liu122footnotemark: 2   Flood Sung2  Xiuying Chen3     Shuo  Shang4   Rui  Yan1,5,6thanks:   Corresponding authors.
+1Gaoling School of Artificial Intelligence, Renmin University of China  2Moonshot AI 3Mohamed bin Zayed University of Artificial Intelligence 4University of Electronic Science and Technology of China  5School of Artifcial Intelligence, Wuhan University 6Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MoE {xiaoqingz, yuhan.liu, ruiyan}@ruc.edu.cnfloodsung@moonshot.cn xy-chen@pku.edu.cnjedi.shang@gmail.com
 1 Introduction
@@
 (2024); Liu et al.
-(2024); Zhang et al.
+(2024a); Zhang et al.
 (2024), and software testing Li et al.
@@
 In the first direction of path optimization, for example, Self-Planning Jiang et al.
-(2024) reduces problem complexity in code generation by introducing a planning phase, which refines the process of solution generation.
+(2024) reduces problem complexity in code generation by introducing a planning phase, which refines the solution generation process.
 Similarly, CodeCoT Huang et al.
-(2023a) improves the exploration by employing a Chain-of-Thought prompt, enabling models to reason through each step systematically.
+(2023a) improves the exploration by employing a Chain-of-Thought(CoT) prompt, enabling models to reason through each step systematically.
 In the second direction, the integration of reflection, test cases, and feedback is progressively being used to steer and enhance the generation process.
@@
 (2023b) boosts efficiency and accuracy by using interactions between multiple agents to provide valuable input.
-Despite their impressive capabilities, a critical challenge remains in fully exploiting the potential of LLMs for high-quality code generation, as their performance is still constrained by issues such as high computation costs and limited effectiveness in integrating iterative feedback.
-To enhance efficiency and minimize trial-and-error costs, it’s essential to define the improvement direction clearly before implementing refinement.
-While existing methods like MapCoder, which leverages four LLM-based agents, consumes a considerable amount of computational resources.
-To address these challenges, we propose ThinkCoder, a novel framework for efficient code generation that combines thorough exploration and optimal refinement.
-First, we simplified our multi-agent framework, which now consists only of the Exploration Agent and the Execution Agent.
-The Exploration Agent is responsible for deep thinking and expanding exploration with various codes and tests.
-The Execution Agent independently verifies code during the optimal refinement process, without relying on LLMs.
-By utilizing a dynamically created testing pool, ThinkCoder ensures accurate verification of the generated code and offers reliable guidance for further optimization.
+Despite the significant potential of using CoT for planning code improvements, the thought process incurs a substantial token overhead.
+To enhance efficiency and minimize trial-and-error costs, it is crucial to clarify the optimization direction before implementing refinement.
+Existing methods, such as MapCoder, which leverages four LLM-based agents, consume a considerable amount of computational resources.
+Based on these insights, we propose ThinkCoder, a novel framework for efficient code generation that combines thorough exploration and optimal refinement.
+First, we simplified the multi-agent framework into a single Exploration Agent and a CodeVerifier.
+The Exploration Agent is responsible for in-depth analysis and generating various codes, while exploring diverse test cases to build a robust testing pool.
+The CodeVerifier, using the testing pool, independently determines the optimization direction without relying on LLMs.
+By continuously improving the testing pool throughout the refinement process, ThinkCoder ensures accurate verification of generated code in each round, providing efficient guidance for subsequent optimizations.
 Second, to tackle the additional computational load resulting from extensive exploration, we incorporate a preference-driven optimization phase using Reinforced Self-Training (ReST) in ThinkCoder.
@@
 Our contributions are as follows:
-$\bullet$ We propose ThinkCoder, an efficiency framework for code generation that incorporates only two agents: an exploration agent and an execution agent that enables thorough exploration and optimal refinement, ensuring high-quality code generation.
-$\bullet$ We introduce preference-driven optimization with ReST, using exploration trajectories from ThinkCoder to enhance the generation and verification capabilities of LLMs, further reducing computational overhead during testing.
+$\bullet$ We introduce ThinkCoder, an efficiency framework for code generation that integrates an LLM-based exploration agent with a non-LLM-based CodeVerifier.
+We utilize a self-evolving testing pool, significantly enhancing the speed of code verification and improvement.
+$\bullet$ We introduce preference-driven optimization with ReST, using exploration trajectories from ThinkCoder to enhance LLMs’ generation and verification capabilities, further reducing computational overhead during testing.
 $\bullet$ We have demonstrated the effectiveness of ThinkCoder on multiple benchmark datasets and across different scales of LLMs, achieving state-of-the-art performance in code generation while optimizing computational efficiency.
@@
 Multi-Agent Code Generation.
-Deriving multiple agents by mimicking the different stages of human programming can effectively improve both the quality and efficiency of code generation.
+Deriving multiple agents by mimicking the different stages of human programming can effectively improve both the quality and efficiency of code generation, as agents based on large language models possess strong reasoning capabilities Liu et al.
+(2024b, c); Zhang et al.
+(2025).
 For instance, AgentCoder Huang et al.
-(2023b) introduces programmer agents, test design agents, and test execution agents, integrating code generation, test case generation, and test feedback into the code generation process.
+(2023b) introduces programmer agents, test design agents, and test execution agents during the code generation process.
 Similarly, MapCoder Islam et al.
@@
 Figure 1: The end-to-end process of ThinkCoder involves $k$ thorough exploration steps followed by $n$ optimal refinement cycles.
-The Exploration Agent generates $k$ codes and tests simultaneously, storing results in a Testing Pool.
-The optimal refinement phase includes self-verification that selects the optimal solution with the Execution Agent and aggregate reflection for the instruction update.
+The Exploration Agent generates $k$ codes and tests simultaneously during self-exploration, storing results in a Testing Pool.
+The optimal refinement phase includes self-verification that selects the optimal solution with the CodeVerifier and aggregate reflection for the instruction update.
 The optimal refinement will be repeated recursively for $n$ cycles, ultimately leading to the final solution.
@@
 Given an instruction $I=[p,\text{prompt}]$ for problem $p$, where prompt is the task description for code generation, the Exploration Agent produces $k$ candidate solutions.
-Each solution is paired with $m$ test cases, and the temperature $t$ of the LLM’s generation is adjusted to promote diversity.
+Each solution is paiblue with $m$ test cases, and the temperature $t$ of the LLM’s generation is adjusted to promote diversity.
 These test cases are then aggregated into a Testing Pool $TP$.
 We define the exploration results as $E=[G,TP]$, where $G=\{g_{i}\mid g_{i}=\text{LLM}(p,t)\text{ for }i\in\{1,2,\dots,k\}\}$, and $TP\in\mathbb{T}^{k\times m}$, where $\mathbb{T}$ represents each test.
-The Execution Agent sequentially executes each candidate solution in $G$ using the test cases from $TP$ to compute pass rates.
+The CodeVerifier sequentially executes each candidate solution in $G$ using the test cases from $TP$ to compute pass rates.
 The code $g_{s}$ with the highest pass rate is selected for further refinement.
 Here we define the pass rate of $g_{i}$ as $r_{g_{i}}=\frac{\text{number of tests passed by }g_{i}}{|TP|}$ and $g_{s}=\arg\max_{i}r_{g_{i}}$.
-Feedback $f$ from failed test case $f_{t}$, along with the current code $g_{s}$, generated test cases $tests$, and problem $p$, is used to update the instructions $I=[p,g_{s},f_{t},f,tests,prompt]$ for the next exploration.
+Feedback $f$ from failed test case $f_{t}$, along with the current code $g_{s}$, generated test cases $tests$, and problem $p$, is used to update the instructions $I=[p,g_{s},f_{t},f,\text{tests},\text{prompt}]$ for the next exploration.
 This iterative process allows the Exploration Agent to generate improved solutions and test cases over $n$ iterations.
 To avoid unnecessary computation on well-solved problems, if $r_{g_{i}}$ is greater than $\theta$, the refinement process is terminated, and the code $g_{s}$ is returned as the final solution.
-Please refer to the ThinkCoder algorithm 1 for more detailed information on these operations.
-Below, we detail the implementation of Thorough Exploration, Optimal Refinement, Exploration&Execution Agent, and Testing Pool.
-Algorithm 1 ThinkCoder
-1:Initialize the testing pool TP=[], set the globally max pass rate $g_{p}=0$ and best solution $g_{s}=None$
-2:Set the hyperparameters $t=0.5,k=5,n=5,m=3,\theta=0.8$
-3:Set the instruction $prompt$ for code generation task and instruction for task $p$ as $I=[p,prompt]$
-4:for each $iter\in n$ do
-5:     Initialize the locally max pass rate $l_{p}=0$ and best solution ${l_{s}}={None}$ for iteration $iter$
-6:     Randomly select $m$ successful test cases from TP as $tests$
-7:     Select a new type of error feedback for test $f_{t}$ from TP as $f$
-8:     if $f_{t}$ and $f$ is not None then
-9:         Update the instruction $I=[p,g_{s},f_{t},f,tests,prompt]$
-10:     end if
-11:     Get code solutions $G^{\prime}$ and test cases $TP^{\prime}$ from exploration agent for the instruction $I$ within $k$ iterations
-12:     Deduplicate testing pool TP’ by AST
-13:     for each $i\in|G^{\prime}|$ do
-14:         Get the pass rate $r_{g_{i}}$ for solution $g_{i}$ from execution agent with the input $TP^{\prime}$
-15:         if $r_{g_{i}}>l_{p}$ then
-16:              Set $l_{p}=r_{g_{i}},l_{s}={g_{i}}$
-17:         end if
-18:     end for
-19:     if $l_{p}>g_{p}$ then
-20:         Set $g_{p}=l_{p},g_{s}=l_{s}$
-21:         Add TP’ to TP
-22:     end if
-23:     if $g_{p}>\theta$ then
-24:         return $g_{s}$ as the best solution
-25:         Terminate optimal refinement iteration
-26:     end if
-27:end for
+For more details of ThinkCoder please refer to the algorithm 1 in Appendix A.1.
+Below, we detail the implementation of Thorough Exploration, Optimal Refinement, Exploration Agent&CodeVerifier, and Testing Pool.
 3.2 Thorough Exploration
-Before running the generated code and performing corrections, ThinkCoder conducted extensive thinking during the Thorough Exploration phase.
-By generating $k$ answers at once, it expanded the search area.
-The search for multiple answers was achieved by increasing the LLM’s temperature.
-With the increasing temperature, the probability distribution of the LLM becomes flatter, reducing the differences between word probabilities and increasing the entropy of the distribution, which enhances the diversity of the generated text.
-The first iteration of exploration aims to generate diverse code and comprehensive test cases, including regular tests, boundary tests, and performance tests.
+During the thorough exploration phase, ThinkCoder conducts extensive thinking before running code and performing corrections.
+The self-exploration process generates $k$ answers simultaneously to expand the search area by increasing the LLM’s temperature, which flattens the probability distribution and enhances diversity.
+The first iteration of self-exploration aims to generate diverse code and comprehensive test cases, including regular tests, boundary tests, and performance tests.
 While the subsequent iterations, focus on the exploration of code correction for specific errors.
-The exploration is carried out by the Exploration Agent, which is the only LLM used in ThinkCoder.
-The prompt design refers to Appendix A.3.
+The only LLM-based agent conducts self-exploration with prompt design detailed in Appendix A.4.
 3.3 Optimal Refinement
@@
 Therefore, optimal refinement involves selecting the best solution via self-verification and gathering the necessary information for deeper exploration and improvement in the next exploration.
-During self-verification, code quality is evaluated by the pass rate, determined by the Execution Agent running candidate code against all test cases.
-The code with the highest pass rate is selected for further refinement.
-From this, we randomly select an unresolved error to target for the current round of improvements.
-This ensures all errors are eventually addressed.
-The refinement instruction is created using the following information: the problem, the best solution identified so far, feedback, and test cases.
+During self-verification, code quality is assessed by the pass rate, calculated by the CodeVerifier running candidate code against all test cases.
+The code with the highest pass rate is chosen for refinement.
+We then randomly select an unresolved error to target in the current round of improvements, ensuring all errors are eventually addressed.
+The refinement instruction is generated based on the problem, the current best solution, feedback, and test cases.
 After each refinement, the solution’s pass rate is evaluated.
-If it surpasses any previous solution, it is marked as the best.
-Through multiple iterations, the best solution is continuously corrected and optimized, leading to a gradual increase in the pass rate within the testing pool.
+If it exceeds previous solutions, it becomes the new best.
+Through multiple iterations, the best solution is continuously corrected and optimized, resulting in a gradual increase in the pass rate within the testing pool.
 For more details, please refer to the Appendix A.4.
-3.4 Exploration & Execution Agent
-ThinkAgent consists of two agents: the exploration agent and the execution agent.
-The exploration agent uses LLMs to search for diversity code and test cases, while the execution agent runs the provided code snippets and test cases in a local environment to generate feedback.
-To improve the efficiency of the execution agent, we leverage multiprocessing to validate the code in parallel.
+3.4 Exploration Agent& CodeVerifier
+ThinkAgent consists of a single exploration agent and a CodeVerifier.
+The exploration agent uses LLMs to search for diversity code and test cases, while the CodeVerifier runs the provided code snippets and test cases in a local environment to generate feedback.
+To improve the efficiency of the CodeVerifier, we leverage multiprocessing to validate the code in parallel.
 3.5 Testing Pool
-The Testing Pool contains the test cases generated by the exploration agent which contains significantly more test cases than the ground truth tests.
-These test cases are categorized into three types: regular testing, boundary testing, and performance testing.
-A higher pass rate in the testing pool reflects stronger code robustness, making it essential to ensure the diversity and accuracy of the test cases.
+The Testing Pool includes numerous test cases generated by the exploration agent, categorized into regular, boundary, and performance tests.
+Our goal is to create a robust pool to assess complex situations beyond simple cases.
+A higher pass rate indicates stronger code robustness, so we ensure test case diversity and accuracy.
 Several key operations contribute to achieving this.
@@
 Finally, the testing pool will select test cases with distinct errors and use feedback to update the code’s improvement goals, driving continuous optimization.
+4 Preference-driven Optimization
+In ThinkCoder, the generated code evolves as the pass rate on the testing pool improves.
+Each increase in the pass rate signifies the development of better codes.
+This progression demonstrates that the exploration agent is gradually producing higher-quality outputs.
+By collecting the trajectory of these improvements, we aim to fine-tune the model, enabling the LLM to generate superior code and test cases more effectively.
+This approach helps to refine the exploration process, clarify its direction, and reduce the time required for exploration.
+Figure 2 describes the process of trajectory collection and fine-tuning of the LLM.
+We will explain this process from the perspectives of Data Collection, the Model Training and Inference.
 […]
@@
 Percentages in parentheses show improvement over the base LLMs.
-4 Preference-driven Optimization
-In ThinkCoder, the generated code evolves as the pass rate on the testing pool improves.
-Each increase in the pass rate signifies the development of better codes.
-This progression demonstrates that the exploration agent is gradually producing higher-quality outputs.
-By collecting the trajectory of these improvements, we aim to fine-tune the model, enabling the LLM to generate superior code and test cases more effectively.
-This approach helps to refine the exploration process, clarify its direction, and reduce the time required for exploration.
-Figure 2 describes the process of trajectory collection and fine-tuning of the LLM.
-We will explain this process from the perspectives of Data Collection, the Model Training and Inference.
 4.1 Data Collection
@@
 We generate high-quality preference data $D_{M_{0}}\cup R_{M_{0}}$ for LLM fine-tuning through the temperature-based and reflection-based generation process.
-Temperature-based Generation.
-During temperature-based generation, we obtain $D_{M_{0}}$ to enhance the diversity of responses.
-For a given problem $x$, the language model $\mathcal{M}_{0}$ is used to explore $k$ times, generating $\{y^{j}\}_{j=1}^{k}$, where $y^{j}\sim\mathcal{M}_{0}(y|x)$.
-During exploration, a relatively large temperature value $t$ is set for $\mathcal{M}_{0}$.
-The generated outputs are then evaluated using the ExecuteAgent, which computes $E(x,y^{j})=(pass\_rate,feedback)$.
-For any sample where $pass\_rate==1$, the corresponding instance $(x,y^{j})$ is added to the training dataset $D_{M_{0}}$.
-Reflection-based Generation.
-The primary purpose of generating $R_{M_{0}}$ is to enhance the ability of error correction.
-For a given sample $x$ and the feedback $E(x,y^{j})=(pass\_rate,feedback)$ from the ExecuteAgent, for samples with $pass\_rate=0$, we add the feedback to $x$ and use the language model $\mathcal{M}_{0}$ to perform temperature-based exploration on the input $x^{\prime}=(x,feedback)$ and get the output $y^{\prime j}$.
-Then we also use the execution agent to obtain the right correction answer with $E(x^{\prime},y^{\prime j})=(pass\_rate^{\prime},feedback^{\prime})$.
-Finally, we add the instance $(x^{\prime},y^{\prime j})$ with the $pass\_rate^{\prime}==1$ to the reflection-based trajectories $R_{M_{0}}$.
+[…]
+[…]
+[…]
+GPT-3.5-Turbo MapCoder 78.3 54.4 80.5 70.1 70.8
+AgentCoder 89.9 89.1 79.9 77.4 84.1
+ThinkCoder(Ours) 90.9 78.8 89.0 81.0 84.9
+GPT-4-Turbo MapCoder 83.1 57.7 93.9 82.9 79.4
+AgentCoder 91.4 91.4 89.6 76.2 87.2
+ThinkCoder(Ours) 90.5 80.5 93.9 86.0 87.7
+Table 2: Pass@1 for MapCoder, AgentCoder, and ThinkCoder on MBPP, MBPP-ET, HumanEval, and HumanEval-ET benchmarks with GPT-3.5-Turbo and GPT-4-Turbo as the backbone.
+[…]
+GPT-4o LDB 82.4 65.4 98.2 81.7 29.3 71.4
+LPW 84.8 65.8 98.2 84.8 34.7 73.7
+ThinkCoder(Ours) 90.0 79.2 97.0 89.6 40.1 79.2
+Table 3: Pass@1 for LDB, LPW, and ThinkCoder on MBPP, MBPP-ET, HumanEval, HumanEval-ET, and CodeContests benchmarks with GPT-4o as the backbone.
 4.2 Model Training and Inference
-Generating high-probability text does not necessarily align well with human preferences, which may lead to biases in various tasks Perez et al.
+Generating high-probability text does not necessarily align well with human preferences, which may lead to biases in various tasks Becker et al.
 (2024).
@@
 We train the base model $\mathcal{M}$ with the success trajectory dataset $D_{M_{0}}\cup R_{M_{0}}$:
-[…]
-In the inference process, two distinct configurations are used.
-The first configuration utilizes the fine-tuned base model, where the input $x$ is fed directly into the model to generate text with a temperature setting of $t=0$.
-This setup is designed to assess whether the data collected through ThinkCoder enhances the quality of the generated output.
-The second configuration regards the fine-tuned base model as the exploration agent within the ThinkCoder framework.
-This approach generates and refines solutions using $k$ explorations and $n$ refinements, with various temperature settings $t$.
-It aims to investigate the framework’s ability to improve the generated solutions through the success trajectories and the ThinkCoder framework.
-Comprehensive experimental results are presented in the evaluation section.
+$\textstyle\mathcal{L}_{MLE}(\theta_{\mathcal{M}})=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathcal{M}_{0}}\cup\mathcal{R}_{\mathcal{M}_{0}}}\log p_{\theta_{\mathcal{M}}}(y\mid x)$
+In inference, we compare the performance with baselines at two configurations.
+The first uses a fine-tuned base model, feeding input $x$ directly to generate text at temperature $t=0$.
+This setup tests if ThinkCoder data improves output quality.
+The second configuration uses the fine-tuned base model as the exploration agent in ThinkCoder.
+It generates and refines solutions with $k$ explorations and $n$ refinements, using different temperature settings $t$.
+This aims to assess the framework’s ability to improve solutions through successful trajectories and the ThinkCoder framework.
 5 Experiments
@@
 Using temperature-based and reflection-based generation methods, we expanded this set to over 2,000 examples for ReST training.
+We also evaluate our framework on two major benchmarks designed for harder problems: LiveCodeBench and CodeContests.
+LiveCodeBench consists of the 167 most recent samples collected between October 2024 and January 2025, while CodeContests includes 165 problems.
 5.2 Baselines
-In terms of LLM base model selection for the application of ThinkCoder, we considered a range of advanced instruction-tuned models, including the general-purpose models LLama2-7B-Chat, Kimi [^1] https://platform.moonshot.cn, GPT-4-Turbo(gpt-4-1106-preview) and GPT-4o, as well as the code-specific model CodeQwen1.5-7B-Chat.
+In terms of LLM base model selection for the application of ThinkCoder, we considered a range of advanced instruction-tuned models, including the general-purpose models LLama2-7B-Chat, Kimi [^2] https://platform.moonshot.cn, GPT-4-Turbo(gpt-4-1106-preview) and GPT-4o(gpt-4o-2024-05-13), as well as the code-specific model CodeQwen1.5-7B-Chat.
 We conducted a comprehensive evaluation of ThinkCoder’s Pass@1 performance on MBPP and HumanEval across LLMs of different parameter sizes.
 For the test-time optimization process, we used the base model LLama2-7B to evaluate performance changes on the MBPP dataset.
-Comparisons were made between results obtained before and after fine-tuning, using either the original training data or the success trajectories, and incorporating the ThinkCoder framework or excluding it.
-For the composite model, we compared it with the current State-of-the-Art (SOTA) models, Self-Collaboration Dong et al.
-(2024), MapCoder Islam et al.
-(2024), and AgentCoder Huang et al.
-(2023b).
-Models Optimal Refinement Iterations
-0 1 2 3 4 5
-LLama2-7B 14.7 22.7 24.9 27.7 32.7 36.9
-LLama2-7B(SFT) 27.8 35.3 37.5 37.6 38.0 38.1
-LLama2-7B(ReST) 32.3 46.2 44.8 46.1 45.7 46.1
-Table 2: Performance comparison of the LLama2-7B model under different settings: baseline without fine-tuning, supervised fine-tuning (SFT) with original train dataset, and Reinforced Self-Training (ReST) with success trajectories from ThinkCoder on the MBPP dataset.
+We compared results obtained before and after fine-tuning, using either the original training data or the success trajectories and incorporating or excluding the ThinkCoder framework.
+For the composite model, we compared it with the current SOTA models, MapCoder Islam et al.
+(2024), AgentCoder Huang et al.
+(2023b), LDB Zhong et al.
+(2024), and LPW Lei et al.
+(2024).
+[…]
+GPT-4-Turbo 0 33.4 16.1
+1 43.7(30.8%) 26.1(62.1%)
+2 46.1(38.0%) 27.3(69.6%)
+5 44.9(34.4%) 28.5(77.0%)
+GPT-4o 0 22.5 21.0
+1 35.3(56.9%) 34.1(62.4%)
+2 38.3(70.2%) 38.9(85.2%)
+5 41.3(83.6%) 40.1(91.0%)
+Table 4: ThinkCoder performance on LiveCodeBench and CodeContests. $n$ represents the self-refinement budget, with the self-exploration budget fixed at $k=5$.
 Figure 3: (a) The Pass@1 metric of the baseline models under different temperature $t$.
@@
 The exploration agent generates $m=3$ test cases per code.
-To compare ThinkCoder with state-of-the-art (SOTA) results, we set the hyperparameters $k$, $t$, $n$, $\theta$, and $m$ to 20, 1.0, 2, 1.0, and 3, respectively.
+To compare ThinkCoder with SOTA results, we set the hyperparameters $k$, $t$, $n$, $\theta$, and $m$ to 20, 1.0, 2, 1.0, and 3, respectively.
 In the data collection phase, the exploration agent $\mathcal{M}_{0}$ is specified as CodeQwen1.5-7B-Chat because it is free and performs well.
@@
 (2021) to train LLMs.
-Figure 4: The variation in Pass@1 performance of the exploration code and test cases on the MBPP dataset under ThinkCoder, as evaluated by CodeQwen1.5-7B-Chat as the base LLM.
 6 Evaluation
 Table 1 illustrates the performance changes of various base LLMs on the MBPP and HumanEval datasets as the Optimal Refinement iteration $n$ increases.
-Table 3 compares the performance of ThinkCoder and SOTA models, using GPT-4-Turbo as the base LLM.
+Table 2 and Table 3 compare the performance of ThinkCoder and SOTA models, using GPT-4-Turbo and GPT-4o as the base LLMs.
+Table 4 shows great potential on harder benchmarks.
+Table 5 compares ThinkCoder’s computation load with MapCoder at similar settings.
 Table 6 presents the performance changes of LLama2-7B before and after fine-tuning with trajectory data collected by ThinkCoder.
+Table 7 illustrates the ablation study of self-exploration and self-refinement.
 Figure 3 shows the impact of changes in the hyperparameter temperature $t$, exploration budget $k$, and refinement iterations $n$ on the Pass@1 and Pass@k metric.
-Figure 4 depicts the trends in the accuracy of generated test cases and generated codes as the optimal refinement iterations in the ThinkCoder increase.
-Table 4 compares ThinkCoder’s computation load with MapCoder at similar settings.
-[…]
-Self-Collaboration 90.2 70.7 78.9 62.1 75.5
-MapCoder 83.1 57.7 93.9 82.9 79.4
-AgentCoder 91.4 91.4 89.6 76.2 87.2
-ThinkCoder 90.5 80.5 93.9 86.0 87.7
-Table 3: End-to-end results of ThinkCoder and state-of-the-art (SOTA) baselines.
+Figure 4 depicts the trends in the accuracy of generated test cases and codes as the optimal refinement iterations in the ThinkCoder increase.
 6.1 Performance on Code Generation
@@
 When the number of exploration iterations $k$ is extended to 20, ThinkCoder surpasses AgentCoder’s performance, which needs 5 refinement rounds, in just 2 iterations.
-Models Exploration k HumanEval MBPP
-Pass@1 Refinement n Tokens (k) Pass@1 Refinement n Tokens (k)
-MapCoder 5 93.9 5 10.41 83.1 3 4.96
-ThinkCoder 5 90.2 1 1.92 89.5 1 1.44
+Models HumanEval Mbpp CodeContest
+n k Pass@1$\uparrow$ Tokens (k)$\downarrow$ n k Pass@1$\uparrow$ Tokens (k)$\downarrow$ n k Pass@1$\uparrow$ Tokens (k)$\downarrow$
+MapCoder 5 5 93.9 21.3 3 3 83.1 5.58 3 3 28.5 18.3
+ThinkCoder(Ours) 20 1 93.9 0.26 5 1 89.5 0.24 5 5 28.5 2.94
 Table 5: The token usage for agent responses and the required refinement iterations in ThinkCoder for HumanEval and MBPP, compared to MapCoder (as ThinkCoder and MapCoder share similar configurations), with GPT-4-Turbo as the base LLM.
 The tiktoken package is used to calculate the response token usage.
+Figure 4: The variation in Pass@1 performance of the exploration code and test cases on the MBPP dataset under ThinkCoder, as evaluated by CodeQwen1.5-7B-Chat as the base LLM.
+6.2 Ablation Study of Self-exploration and Self-refinement
+Figures 3(b) and 3(c) highlight the performance gains from both self-exploration and self-refinement.
+The gain from self-exploration is greater than that from self-refinement.
+As illustrated in Table 7, self-refinement significantly enhances efficiency when paired with self-exploration.
+However, increasing the exploration budget excessively (e.g., $k=20$) leads to inefficiency.
+A moderate increase in the exploration budget accelerates the optimization of the self-refinement process.
+Models Optimal Refinement Iterations
+0 1 2 3 4 5
+LLama2-7B 14.7 22.7 24.9 27.7 32.7 36.9
+LLama2-7B(SFT) 27.8 35.3 37.5 37.6 38.0 38.1
+LLama2-7B(ReST) 32.3 46.2 44.8 46.1 45.7 46.1
+Table 6: Performance comparison of the LLama2-7B model under different settings: baseline without fine-tuning, supervised fine-tuning (SFT) with original train dataset, and Reinforced Self-Training (ReST) with success trajectories from ThinkCoder on the MBPP dataset.
 6.3 Performance with ReST Training
@@
 Secondly, with a temperature setting of 0.5, the model achieves optimal performance with only a single round of refinement, underscoring the efficiency of ReST fine-tuning in reducing computational overhead.
+Self-exploration Budget k Self-refinement Budget n
+1 2 3 5 10 20
+1 91.7 90.0 91.7 93.3 93.3 95.0
+2 93.3 95.0 95.0 96.7 96.7 96.7
+5 93.3 93.3 95.0 95.0 96.7 96.7
+10 96.7 96.7 96.7 96.7 96.7 96.7
+20 96.7 96.7 96.7 96.7 96.7 96.7
+Table 7: The performance trends on the randomly 60 Mbpp samples on GPT-4 with varying compositions of self-exploration and self-refinement budgets.
 6.4 Evolution of Generated Code and Tests
@@
 As shown in Figure 3(a), models exhibit optimal performance at different temperature settings.
-LLama2-7B-Chat, CodeQwen1.5-7B-Chat and Kimi improve in Pass@1 as temperature increases, while GPT-4-Turbo and GPT-4o perform better at lower temperatures.
-As depicted in Figure 3(b), the Pass@k performance of various base LLMs improves with an increasing exploration budget $k$, highlighting the models’ ability to find optimal solutions.
-However, the performance eventually levels off, suggesting that the Pass@k performance of these models will eventually stabilize.
-As shown in Figure 3(c), Pass@1 performance of base LLMs improves with increasing refinement iterations $n$, following a trend similar to the budget curve during exploration.
-However, exploration consistently achieves greater performance gains, underscoring the importance of optimizing the LLM’s answer selection ability for better code quality.
-For the hyperparameter study of $\theta$ and $m$, please refer to Appendix A.1.
+LLama2-7B-Chat, CodeQwen1.5-7B-Chat, and Kimi improve in Pass@1 as temperature increases, while GPT-4-Turbo and GPT-4o perform better at lower temperatures.
+Figure 3(b) indicates that Pass@k performance of base LLMs improves with increasing exploration budget $k$ but eventually stabilizes.
+Figure 3(c) shows that Pass@1 performance improves with more refinement iterations $n$, but exploration gains are more significant.
+For details on hyperparameters $\theta$ and $m$, see Appendix A.2 and A.3.
 6.6 Cost Analysis
-In Figure 4, we compare the computational overhead between MapCoder and ThinkCoder.
-On the HumanEval and MBPP datasets, a single refinement with ThinkCoder outperforms the result of five iterations of refinement with MapCoder, while using fewer tokens.
-This is because MapCoder utilizes four agents that are all powered by LLMs, whereas our framework uses a single LLM to handle both code and test generation, while debugging is directly based on feedback from the Python environment, significantly reducing token usage.
+In Table 5, we compare the computational overhead between MapCoder and ThinkCoder.
+Compablue to MapCoder, which uses four LLM-based agents executed serially, ThinkCoder employs only one LLM-based agent and one non-LLM-based CodeVerifier, resulting in lower time overhead.
+Therefore, we use token overload as a measure of cost consumption.
+It reveals that ThinkCoder demonstrates more efficient resource utilization compared to MapCoder with a similar pass@1.
 7 Conclusion
-In this paper, we introduce ThinkCoder, a framework designed to improve the code generation capabilities of LLMs.
-The proposed approach incorporates an extensive exploration to generate diverse solutions, followed by an optimal refinement stage that integrates the solutions, tests, and feedback as reflections to enhance performance.
-To minimize computational costs during exploration, we capture successful trajectories through ThinkCoder, which are then leveraged to fine-tune the LLM.
-This process not only boosts the model’s performance in code generation but also reduces the number of refinement iterations needed to achieve optimal solutions.
+In this paper, we introduce ThinkCoder, a framework designed to improve the code generation efficiency of LLMs.
+The proposed approach incorporates LLM’s exploration with a non-LLM-based CodeVerifier that greatly reduces the token usage.
+To further minimize computational costs during exploration, we capture successful trajectories through ThinkeCoder to fine-tune the LLM.
+This process enhances code generation performance while reducing refinement iterations for optimal solutions.
 We evaluate our framework across LLMs of varying sizes and demonstrate its superior performance over current composite models.
+Acknowledgments
+This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building worldclass universities (disciplines) of Renmin University of China.
 Limitations
 In our current approach to managing the testing pool, we primarily focus on selective expansion but lack effective strategies for cleaning noise from the test set.
-Moreover, the diversity of the test set heavily depends on the generation capabilities of LLMs.
-Our work highlights the importance of tests and identifies the gap between generated and annotated cases, so we plan to address the enhancement of the testing pool in future work.
+Moreover, test set diversity relies heavily on LLMs’ generation capabilities.
+Our work highlights the importance of tests and the gap between generated and annotated cases.
+We plan to enhance the testing pool in future work.
 Furthermore, while ReST training has demonstrated excellent results in enhancing the code generation capabilities of foundational LLMs, validation in ThinkCoder reveals rapid convergence.
 This indicates that fine-tuning teaches the LLM to find good answers, but the improvement in exploring the overall solution space remains limited.
-To address this, we look forward to introducing on-policy training methods in future work, combining data collection with fine-tuning to enable the model to evolve continuously during training, generating more diverse answers and further boosting the LLM’s online exploration capabilities.
+To address this, we look forward to introducing on-policy training methods in future work, combining data collection with fine-tuning to continuously evolve the model during training.
+This will generate more diverse answers and enhance the LLM’s online exploration capabilities.
 Appendix A Appendix
-A.1 Exploration Budget Control
+A.1 Algorithm of ThinkCoder
+Algorithm 1 shows the pseudo-code of ThinkCoder.
+A.2 Exploration Budget Control
 Figure 5: The relationship between the computational cost and performance of ThinkCoder at different budget control thresholds, where ‘Pass@1’ indicates the performance on the MBPP dataset, and ‘$\theta$’ refers to the budget control threshold that allows the task to execute the next exploration process.
@@
 In our experiments, we set $\theta=0.8$, as this value not only progressively improves Pass@1 but also reduces the number of tasks per iteration by nearly half.
-A.2 Exploration for Tests
+Algorithm 1 ThinkCoder
+1:Initialize the testing pool TP=[], set the globally max pass rate $g_{p}=0$ and best solution $g_{s}=\text{None}$
+2:Set the hyperparameters $t=0.5,k=5,n=5,m=3,\theta=0.8$
+3:Set the instruction prompt for code generation task and instruction for task $p$ as $I=[p,\text{prompt}]$
+4:for each $iter\in n$ do
+5:     Initialize the locally max pass rate $l_{p}=0$ and best solution ${l_{s}}={\text{None}}$ for iteration $iter$
+6:     Randomly select $m$ successful test cases from TP as tests
+7:     Select a new type of error feedback for test $f_{t}$ from TP as $f$
+8:     if $f_{t}$ and $f$ is not None then
+9:         Update the instruction $I=[p,g_{s},f_{t},f,\text{tests},\text{prompt}]$
+10:     end if
+11:     Get code solutions $\text{G}^{\prime}$ and test cases $\text{TP}^{\prime}$ from exploration agent for the instruction $I$ within $k$ iterations
+12:     Deduplicate testing pool TP’ by AST
+13:     for each $i\in|\text{G}^{\prime}|$ do
+14:         Get the pass rate $r_{g_{i}}$ for solution $g_{i}$ from CodeVerifier with the input $\text{TP}^{\prime}$
+15:         if $r_{g_{i}}>l_{p}$ then
+16:              Set $l_{p}=r_{g_{i}},l_{s}={g_{i}}$
+17:         end if
+18:     end for
+19:     if $l_{p}>g_{p}$ then
+20:         Set $g_{p}=l_{p},g_{s}=l_{s}$
+21:         Add TP’ to TP
+22:     end if
+23:     if $g_{p}>\theta$ then
+24:         return $g_{s}$ as the best solution
+25:         Terminate optimal refinement iteration
+26:     end if
+27:end for
+A.3 Exploration for Tests
 Metrics Tests m/Exploration
@@
 Overload(s) 28.7 86.0 112.3 364.5
-Table 5: The variation in the Pass@1 metric and the execution agent’s overload is influenced by changes in the number of test cases $m$ during each exploration.
-We randomly selected 50 tasks from the MBPP test data to observe the performance improvement and code execution overhead as the number of test cases $m$ increases during each exploration.
+Table 8: The variation in the Pass@1 metric and the CodeVerifier’s overload is influenced by changes in the number of test cases $m$ during each exploration.
+We randomly selected 60 tasks from the MBPP test data to observe the performance improvement and code execution overhead as the number of test cases $m$ increases during each exploration.
 Table 8 illustrates this trend.
@@
 We observed that increasing $m$ leads to improved Pass@1 performance and results in longer execution times due to the additional effort required for self-verification.
-To balance time efficiency and performance, we set $m$ for all experiments.
-A.3 Prompts for Exploration Agent
+We set $m$ for all experiments to balance time efficiency and performance.
+A.4 Prompts for Exploration Agent
 Figure 6: ThinkCoder exploration agent prompt example.
 Figure 6 illustrates the input received by the exploration agent, including the problem, the best solution, the test cases from the ground truth, the generated test cases from testing pooling, and feedback on test failures.
-A.4 Prompts for Execution Agent
-Figure 7: ThinkCoder execution agent prompt example.
+A.5 Inputs for CodeVerifier
+Figure 7: ThinkCoder CodeVerifier input example.
 Figure 7 shows the input of the executable Python environment, including the code and all test cases from the testing pool.
-A.5 Example of Thorough Exploration
+A.6 Example of Thorough Exploration
 Figure 8: ThinkCoder exploration examples.
 Figure 8 illustrates multiple solutions explored by the exploration agent.
-A.6 Example of Optimal Refinement
+A.7 Example of Optimal Refinement
 Figure 9: ThinkCoder refinement examples.
@@
 The success rate improved from 0.217 to 0.356, leading us to select the current solution as the best solution.
-A.7 Example of Testing Pool
+A.8 Example of Testing Pool
 Figure 10: ThinkCoder testing pool examples.